Water is flowing in a fire

hose with a velocity of 1.0 m/s and a pressure of 200000 Pa. At the nozzle the

pressure decreases to atmospheric pressure (101300 Pa), there is no change in

height. Use the Bernoulli equation to calculate the velocity of the water

exiting the nozzle. (Hint: The density of water is 1000 kg/m3 and

gravity g is 9.8 m/s2. Pay

attention to units!)]


1
Expert's answer
2022-01-25T06:51:37-0500

The output velocity of the water, calculated by Bernoulli's principle, is


"P_1+\dfrac12\rho v_1^2=P_2+\dfrac12\rho v_2^2,\\\space\\
v_2=\sqrt{\dfrac{2(P_1-P_2)}{\rho}+v_1^2}=14\text{ m/s}."